OpenStudy (anonymous):

3(y+6) = 30?? solve the equation

OpenStudy (anonymous):

1) divide by 3 2) subtract 6

OpenStudy (anonymous):

3y +18 = 30 3y = 12 y = 4

OpenStudy (anonymous):

OR like satellite 3(y+6) = 30 y+6 = 10 y = 4